determine the [oh] of a solution with a PH of 5.40

PRACTICE – DETERMINE THE [OH-] OF A SOLUTION WITH A PH OF 5.40
Answer #1

Given : pH = 5.40

We have relation, pH = - log [ H + ]

\therefore 5.40 = - log [ H + ]

Taking anti log on both sides , we get

[ H + ] = 10 -5.40 = 3.98 \times 10 -06 M

We have relation, [ H + ] [ OH - ] = K w = 10 -14

\therefore  [ OH - ] = 10 -14 / [ H + ] = 10 -14 / 3.98 \times 10 -06 = 2.51 \times 10 -09 M

ANSWER : [ OH - ] = 2.51 \times 10 -09 M